OpenStudy (anonymous):

if 3-3i and z2=7(cos(5pi/9)+isin(5pi/9)) then z1/z2=?? PLEASE HELP :)

OpenStudy (anonymous):

A) 21(cos(25pi/36)+i sin(25pi/36)) B) 21sqrt2(cos(29pi/36)+i sin(29pi/36)) C) 3sqrt2/7(cos(29pi/36)+i sin(29pi/36)) D) 3sqrt2/7(cos(-11pi/36)+i sin(-11pi/36))

OpenStudy (kropot72):

\[\frac{z _{1}}{z _{2}}=\frac{r _{1}e ^{i \theta _{1}}}{r _{2}e ^{i\ \theta _{2}}}=\frac{r _{1}}{r _{2}}e ^{i(\theta _{1}}-\theta _{2})\]

OpenStudy (anonymous):

which one of the options do you think it is?

OpenStudy (kropot72):

You need to put z1 and z2 into exponential form and then use the equation above. z1 has to be converted to polar form. \[r=\sqrt{(3^{2}}+3^{2})=\sqrt{18}\] \[\tan \theta=\frac{-1}{1}=-1\] Can you now put z1 into the form\[z _{1}=r \cos \theta+ir \sin \theta\]

OpenStudy (anonymous):

I did, I got D as an answer but I'm not sure it's correct..

OpenStudy (anonymous):

do you agree on D?

OpenStudy (anonymous):

@kropot72

OpenStudy (kropot72):

No. D is not the correct choice.

OpenStudy (anonymous):

well then I did something wrong. lol

OpenStudy (anonymous):

C?

OpenStudy (kropot72):

I already gave you what is needed for 'r' of the answer. \[r=\frac{z _{1}}{z _{2}}=\frac{\sqrt{18}}{7}=\frac{3\sqrt{2}}{7}\] So the only remaining choice is obvious.

OpenStudy (anonymous):

actually the correct answer was D. -.-